A 0.140-kg baseball moves towards home plate with a Velocity initial = -42.0 m/s x-hat. After striking the bat, the ball moves with a Velocity final = 36.0 m/s x-hat + 18.0 m/s y-hat. Find the magnitude of the impulse delivered to the ball by the bat.
Answer #1

here,

mass of baseball , m = 0.14 kg

initial velocity , u = - 42 i m/s

the final velocity , v = 36 i m/s + 18 j m/s

the impulse delevired , I = m * ( v - u)

I = 0.14 * ( 36 i + 18 j - (-42 i)) kg.m/s

I = ( 10.92 i + 2.52 j ) kg.m/s

the magnitude of the impulse delevired to the ball by the bat , |I| = sqrt(10.92^2 + 2.52^2)

|I| = 11.2 kg.m/s